OpenStudy (anonymous):

Please Help? If \(\ P_1=2.00*10^5 \), solve these simultaneous equations for \(\ P_2 \): \(\ T_2=3T_1 \), \(\ P_1V_1^\frac{5}{3} = P_2V_2^\frac{5}{3} \), \(\ \frac{P_1V_1}{T_1} = \frac{P_2V_2}{T_2} \)?

OpenStudy (anonymous):

If \(\ \Large P_1=2.00*10^5 \text{solve these simultaneous equations for } P_2 \): \(\ T_2=3T_1 \), \(\ P_1V_1^\frac{5}{3} = P_2V_2^\frac{5}{3} \), \(\ \frac{P_1V_1}{T_1} = \frac{P_2V_2}{T_2} \text{?}\)

OpenStudy (anonymous):

\(\ P_1=2.00*10^5 \), E1)\(\ P_2 \): \(\ T_2=3T_1 \), E2)\(\ P_1V_1^\frac{5}{3} = P_2V_2^\frac{5}{3}\) E3)\(\frac {P_1V_1}{T_1} = \frac {P_2V_2}{T_2}\) ..... E1) \(\frac{T_2}{T_1}=3\) E2) \(\frac {P_2}{P1} = (\frac{V_1}{V_2})^\frac{5}{3}\) \((\frac {P_2}{P_1})^\frac{3}{5} = \frac{V_1}{V_2}\) E3) \(\frac{V_1}{V_2}\frac{T_2}{T_1} = \frac {P_2}{P_1}\) Put (E1) and (E2) into E(3) \(3(\frac {P_2}{P_1})^\frac{3}{5} = \frac {P_2}{P_1}\) Now solve for \(P_2\) normally since you already know \(P_1\)

OpenStudy (anonymous):

@NicoMiko \(\ \Huge \text{Thanks!!!} \)